# A Hybrid Anyon-Otto thermal machine

### v1

@@ Abstract
Since the HAO cycle operates like a regular Otto cycle in terms of $\overline{W}$ and $\overline{Q}$, it now becomes meaningful to define the efficiency $\eta=\overline{W}/\overline{Q}_{A(B)}$ for $T_{A(B)}>T_{B(A)}$, provided the cycle is operating in the engine mode. Examining the non-interacting situation with $U=0$, $J_{1}=2.0$ and $J_{2}=1.0$, it is evident from Figs. 6-6 that the inverse accelerator mode does not emerge even for $\theta_{1}\neq 0$. However, the effect of changing the statistical properties with $\theta_{1}$ still manifests itself at low temperatures. In particular, for $\theta_{1}\geq 2.2$, the cycle is able to operate as an accelerator for certain range of small but finite temperatures for $T_{A}>T_{B}$. Likewise, for $T_{A}<T_{B}$, the cycle switches from accelerator to engine mode $\theta_{1}\geq 2.2$ at small temperatures. Finally, we also note the emergence of heater mode in the vicinity of the temperatures at which the direction of heat flow switches for the two baths.
@@
Conclusions
@@
In summary, we have proposed a four-stroke hybrid anyon Otto cycle based on the 1D anyon Hubbard model which, like its Pauli counterpart, relies on exclusion statistics of anyons to derive work at low temperatures. In the absence of an explicit interaction among the anyons, we see a monotonous increase in the low temperature work output as the statistical parameter is increased from the bosonic limit to the pseudo-fermionic limit. The presence of a finite anyon energy at low temperature leads to the emergence of an inverse accelerator mode of the cycle, which is prohibited by the second law in a regular Otto cycle. When finite but weak interactions are introduced, and at half or higher filling, the low-temperature work output no longer peaks at the bosonic or pseudo-fermionic limits. Instead, it reaches a maximum at an intermediate value of the statistical parameter, thus, demonstrating that in the interacting regime, anyonic statistics can be harnessed to achieve greater work extraction than is possible with either bosonic or fermionic statistics alone. We reiterate that the inverse accelerator mode does not violate the second law in the HAO cycle; they arise from treating the anyonization stroke - which can be interpreted as a work stroke - as a “heat” source, as clarified in the final discussion. By properly incorporating the anyonic contribution and redefining the work, we show that at low temperatures the accelerator mode is replaced by an engine mode. This engine mode exhibits maximum efficiency in the anyonic limit at high $\theta$, highlighting the crucial role of anyonic statistics in enhancing cycle performance. Finally, we emphasize that the 1D anyon Hubbard model considered in this work has already been realized experimentally and thus the results presented in this work are likely to be experimentally verified in the near future.
@@ Acknowledgements.
ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR"; (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project. R.W.C acknowledges support from the Polish National Science Centre (NCN) under the Maestro Grant No. DEC- 2019/34/A/ST2/00081. U.B. is also grateful for the financial support of the IBM Quantum Researcher Program. T.G. acknowledges the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government, as well as funding by the Department of Education of the Basque Government through the project PIBA_2023_1_0021 (TENINT), and by the Agencia Estatal de Investigación (AEI) through Proyectos de Generación de Conocimiento PID2022-142308NA-I00 (EXQUSMI). A.P. and M.L.B. acknowledge for support from the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+, the Severo Ochoa Grant CEX2023-001292-S, Generalitat Valenciana grant CIPROM/2022/66, the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). M.A G-M acknowledges support from the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call—Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2026 Agenda: also from Projects of MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat Valenciana, with reference 20220883 (PerovsQuTe) and COMCUANTICA/007 (QuanTwin).

### v2

@@ II Results
Since the HAO cycle operates like a regular Otto cycle in terms of $\overline{W}$ and $\overline{Q}$, it now also becomes meaningful to define the efficiency $\eta=\overline{W}/\overline{Q}_{A(B)}$ for $T_{A(B)}>T_{B(A)}$, provided the cycle is operating in the engine mode. Examining the non-interacting situation with $U=0$, $J_{1}=2.0$ and $J_{2}=1.0$, it is evident from Figs. 6-6 that the inverse accelerator mode does not emerge even for $\theta_{1}\neq 0$. However, the effect of changing the statistical properties with $\theta_{1}$ still manifests itself at low temperatures. In particular, for $\theta_{1}\geq 2.2$, the cycle is able to operate as an accelerator for certain range of small but finite temperatures for $T_{A}>T_{B}$. Likewise, for $T_{A}<T_{B}$, the cycle switches from accelerator to engine mode $\theta_{1}\geq 2.2$ at small temperatures. Finally, we also note the emergence of heater mode in the vicinity of the temperatures at which the direction of heat flow switches for the two baths.
@@
Experimental realization
The 1D AHM has already been realized experimentally using ultracold $^{87}\textrm{Rb}$ atoms in an optical lattice [greinerExpt24], where the tunneling amplitude is engineered to depend on the particle density in accordance with Eq. (1). In this section, we propose how the HAO thermal machine can be realized using the same experimental setup. The main challenges associated with implementing the HAO cycle on the optical lattice platform of [greinerExpt24] are the following - (i) realization of a thermal bath with which the system interacts during the thermalization strokes, (ii) measurement of the energy change of the system at the end of each stroke, and (iii) requirement of higher particle density ($N>2$) to observe the superior performance of the cycle for $0<\theta<\pi$.
To overcome the first challenge, we propose to introduce a second set of atoms in the lattice, which acts as a bath for the atoms already present in the system. Specifically, the additional atoms that form the bath will be prepared in the superfluid state ($\theta_{\rm bath}\to 0$) to facilitate a controlled tunability of the effective temperature of the bath. Similar realization of controlled bath environments has been proposed and experimentally implemented in a number of other works [Seetharam_2015, Rubio_Abadal_2019, L_onard_2023, Schnell_2024], and allows for the controlled engineering of dissipation. In our setup, the bath will be located along the length $L$ of the system, where it is prepared in the superfluid state using the Floquet scheme described in [Cardarelli_2016]. The size of the bath will be much larger than the system such that the system can couple to continuum modes. Independent control of the statistical parameter $\theta$ will be necessary for the bath and the system, as $\theta_{\rm bath}\to 0$ while $\theta_{\rm system}$ changes throughout the HAO cycle, and can be realized by implementing the Floquet scheme using a digital micromirror device (DMD) rather than directly modulating the intensity of the lattice. Using the DMD to modulate the intensity allows for spatial control of the statistical parameter, as the DMD can be programmed such that one section projects modulating light to realize the system while another realizes the bath. The system will be separated from the bath by an optically projected wall, implemented either by the same or another DMD, where the height of the wall is determined by the intensity of the light. By changing the intensity of the light, and therefore the height of the wall, the system can be coupled or isolated from the bath during the thermalization or work strokes, respectively.
Concerning the second challenge, we shall resort to time-of-flight (TOF) measurements to find the energy $E(\lambda,\theta)$ at the end of each stroke. We illustrate the measurement protocol only for the non-interacting case ($U=0$) here, as the same protocol will also work for weak interactions. In the absence of interactions, we recall that the work strokes consist of adiabatically tuning the parameter $J$ as the system evolves unitarily, while the heat strokes correspond to a dissipative evolution of the system in contact with thermal baths along with a simultaneous tuning of the phase parameter $\theta$. The measurement protocol will therefore be as follows. Let us assume that the system is initially prepared in the state $\rho_{3}$, i.e., in thermal equilibrium with the bath with temperature $T_{B}$ and $\theta=0$. We first perform a TOF measurement to determine the energy $E(J_{2},0)$ of the system in this state. The system is confined by optically projected walls along both $x$ and $y$, with the walls along $y$ defining the length $L$, and to perform TOF, the walls along $y$ are quenched to allow the system to expand within a 1D tube. Fluorescence images of the atoms are taken at successively longer times after expansion to track the expansion profile of the system, from which a velocity and therefore the energy may be extracted. Since the TOF measurement destroys the initial state, we reset the system to $\rho_{3}$ before starting the HAO cycle. The compression work stroke is performed by decoupling the bath and adiabatically tuning the hopping amplitude $J_{2}\to J_{1}$ over a time-interval $\tau_{W}$, during which the system evolves to $\rho_{4}$. At $t=\tau_{W}$, the TOF measurement is performed again to find the energy $E(J_{1},0)$. The work $W_{21}$ is therefore obtained as $W_{21}=\langle E(J_{1},0)-E(J_{2},0)\rangle$.
In the next step, we once again prepare the system in state $\rho_{3}$, evolve unitarily to $\rho_{2}$, followed by coupling the system to bath with temperature $T_{A}$. The phase parameter is then tuned to $\theta=\theta_{1}$ while in contact with the bath, and we ensure that the system thermalizes with the bath for sufficiently long time after tuning of the phase parameter is completed, before performing the TOF measurement to find $E(J_{1},\theta_{1})$. The heat exchanged is thus calculated as $Q_{A}=\langle E(J_{1},\theta_{1})-E(J_{2},0)\rangle-W_{21}$. Similarly, to calculate $W_{12}$ and $Q_{B}$, we shall reset the system to $\rho_{3}$ in each case, and then perform TOF measurements at the end of the unitary expansion and thermalization B strokes, respectively, to find $W_{12}=\langle E(J_{2},\theta_{1})-E(J_{2},0)\rangle-W_{21}-Q_{A}$ and $Q_{B}=\langle E(J_{2},0)^{\prime}-E(J_{2},0)\rangle-W_{21}-Q_{A}-W_{12}$, where $E(J_{2},0)^{\prime}$ is the measured energy of the system after completing the cycle. We note that $E(J_{2},0)^{\prime}\neq E(J_{2},0)$ if the thermalization is not perfect in the heat exchange strokes, which will indicate an incomplete cycle.
Finally, regarding the third challenge, achieving a higher particle density ($N>2$) is possible as long as the filling $n=N/L\lesssim 0.36$ [Str_ter_2016]; otherwise, the Floquet implementation does not faithfully realize the AHM. When exploring behavior at higher filling, it may be possible that the essential physics is preserved despite differences between the driven system and the AHM. Other implementations that do not involve shaking may be considered to realize systems at higher filling [Fr_lian_2022].
III Methods
All the numerical simulations were performed in the bosonic representation of the AHM Hamiltonian (Eq. (1)) using the QuSpin package [quspin1, quspin2]. The density dependent phase of the hopping terms cannot be directly implemented using the standard Quspin routines for constructing Hamiltonians. Hence, we explicitly encoded the hopping terms as sparse matrices and added the same to the onsite interaction terms constructed directly using QuSpin routines [code_ahm_hamiltonian]. The full Hamiltonian was then diagonalized within Quspin for generation of required data.
IV Discussion
@@
In summary, we have proposed a four-stroke hybrid anyon Otto cycle based on the 1D anyon Hubbard model which, like its Pauli counterpart, relies on exclusion statistics of anyons to derive work at low temperatures. In the absence of an explicit interaction among the anyons, we see a monotonous increase in the low temperature work output as the statistical parameter is increased from the bosonic limit to the pseudo-fermionic limit. The presence of a finite anyon energy at low temperature leads to the emergence of an inverse accelerator mode of the cycle, which is prohibited by the second law in a regular Otto cycle. When finite but weak interactions are introduced, and at half or higher filling, the low-temperature work output no longer peaks at the bosonic or pseudo-fermionic limits. Instead, it reaches a maximum at an intermediate value of the statistical parameter, thus demonstrating that in the interacting regime, anyonic statistics can be harnessed to achieve greater work extraction than is possible with either bosonic or fermionic statistics alone. We reiterate that the inverse accelerator mode does not violate the second law in the HAO cycle; they arise from treating the anyonization stroke - which can be interpreted as a work stroke - as a “heat” source, as clarified in the final discussion. By properly incorporating the anyonic contribution and redefining the work, we show that at low temperatures the accelerator mode is replaced by an engine mode. This engine mode exhibits maximum efficiency in the anyonic limit at high $\theta$, highlighting the crucial role of anyonic statistics in enhancing cycle performance. Finally, we outline an experimental protocol to verify our results using a similar experimental platform that was used to realize the AHM.
Authors contribution
SB and UB conceived the project. MB led the numerical calculations, with additional contributions from SB. SB and UB performed the main analysis and interpretation of the results, with further contributions from ML, TG, and RC. AP and MAG-M also contributed to a part of the analysis and interpretation. JK developed the experimental proposal. All authors contributed to preparation of the manuscript.
@@ Acknowledgements.
ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR"; (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project. R.W.C acknowledges support from the Polish National Science Centre (NCN) under the Maestro Grant No. DEC- 2019/34/A/ST2/00081. U.B. is also grateful for the financial support of the IBM Quantum Researcher Program. T.G. acknowledges the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government, as well as funding by the Department of Education of the Basque Government through the project PIBA_2023_1_0021 (TENINT), and by the Agencia Estatal de Investigación (AEI) through Proyectos de Generación de Conocimiento PID2022-142308NA-I00 (EXQUSMI). A.P. and M.L.B. acknowledge for support from the project PID2023-152724NA-I00, with funding from MCIU/AEI/10.13039/501100011033 and FSE+, the Severo Ochoa Grant CEX2023-001292-S, Generalitat Valenciana grant CIPROM/2022/66, the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). Also funding from Horizon Europe EU projects MSCA-SE CaLIGOLA, Project ID: 101086123, and MSCA-DN CaLiForNIA, Project ID: 101119552. M.A G-M acknowledges support from the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call—Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan—NextGenerationEU within the framework of the Digital Spain 2026 Agenda: also from Projects of MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat Valenciana, with reference 20220883 (PerovsQuTe) and COMCUANTICA/007 (QuanTwin).
Data Availability
All data generated or analysed during this study are included in this article.
Code Availability
The code for constructing the AHM Hamiltonian can be accessed via the link https://doi.org/10.5281/zenodo.17873350.
Competing Interests
All authors declare no financial or non-financial competing interests.